Sick-leave time used by employees of a firm in one month approximates a normal distribution with a mean of 165 hours and a variance of 225. In planning schedules for next month, how much time should be budgeted for sick leave if that amount is to be exceeded with a probability of 0.992?



(Please round your answer to four decimal places if necessary. Thank you!)

Answer by Boreal(15235)

Probability of 0.992 is with a z-value of +2.41
mean is 165
sd is 15, sqrt(V)
as question is written, want MORE than the value found. That means the amount of hours is at the z value of -2.41.
z<=(x-mean)/sd
-2.41=(x-165)/15
-36.15=x-165
x=128.85 hours
That has a 99.2% probability of being exceeded.